OpenStudy (anonymous):

Suppose that f(x)=ln(5+x^2). Use interval notation to indicate where f(x) is concave up and concave down.

OpenStudy (dumbcow):

2nd derivative determines concavity f''=0 inflection point f'' >0 concave up f'' <0 concave down \[f'(x) = \frac{2x}{x^{2}+5}\] \[f''(x) = \frac{10 -2x^{2}}{(x^{2}+5)^{2}}\] Inflection points: \[x = \pm \sqrt{5}\] \[f''(0) = \frac{10}{25} > 0 \] thus from (-sqrt5, sqrt5) f(x) is concave up from (-inf,-sqrt5] and [sqrt5,inf) f(x) is concave down

OpenStudy (anonymous):

Thanks :]

OpenStudy (dumbcow):

your welcome

OpenStudy (dumbcow):

do you get how i got the derivatives?

OpenStudy (anonymous):

yeah, second derivative of f(x). I was ignoring the inflection points :/

OpenStudy (dumbcow):

yeah inflection points tell you when concavity changes